the lenght of a rectangle is 4 feet less than twice the width. the perimeter is 28 feet. find the length and width fo the rectangle.
Answer by vleith(2983)

Perimeter is the sum of all the sides.
Let w be the width.
L = 2w-4
P = w + l + w + l
28 = w + 2w -4 + w + 2w -4
28 = 6w - 8
36 = 6w
6 = w
l = 12 - 4
l = 8
Does that work?
28 = 12 + 16 check!